Lets begin with an efficiency analysis of a gasoline engine. The engine in a large truck takes in 2500 J of heat and deliverTH 2500 J W-500 J Qe=? TcA gasoline engine with a thermal efficiency of 39 % has a power output of 50 hp. How much heat must be supplied per second to
Answer #1

(a) Efficiency =n= 500 1 ) 3/05 TH H 25ou > n = 0-2 b) Also, ME I-9c -0-2-1- SH ge = 1-0-2 2500 2500 2000 I → Ic=0.8 x 2500 HHope it helps.

Kindly like the answer and comment your doubts in the comment section below.